Let g(x) be the transformation of f(x)= x2. Write the rule for g(x) using the change described.

1. horizontal shift 6 units to the left, vertical shift down by 5 followed by a compression horizontally by a factor of 1/3. ______________
1
Expert's answer
2021-09-16T00:46:54-0400

"f(x)=x^2\\y=x^2"

now according to given condition

"y-5=\frac{1}{3}\times(x-6)^2\\y-5=\frac{1}{3}\times(x^2+36-12x)\\y-5=\frac{x^2}{3}+12-4x"

"y=\frac{x^2}{3}+17-4x\\f(x)=\frac{x^2}{3}+17-4x"